# LayerDAG: A Layerwise Autoregressive Diffusion Model for Directed Acyclic Graph Generation

### v1

@@ 1 Introduction
A Directed Acyclic Graph (DAG) is a data structure that represents the order of elements (Bang-Jensen & Gutin, 2008). Unlike linear sequences/chains, which follow a single, straight path from the first to the last element, DAGs incorporate branching and merging, allowing for the modeling of complex dependencies and hierarchies. This flexibility makes DAGs ideal for representing diverse problem domains such as workload behavior during system execution Sridharan et al. (2023), operator dependence for program analysis (Phothilimthana et al., 2023b; Luo et al., 2021; Cortez et al., 2017), dataflows in circuits (Dong et al., 2023), task dependencies in project management (Skiena, 2008; Borenstein, 2000), and cause-effect relationships (Pearl, 1995; Tennant et al., 2020). Additionally, DAGs have recently been used to create challenging benchmarks for evaluating the reasoning capabilities of large language models (Zhu et al., 2024).
@@
Despite the benefits of DAGs mentioned above, it is non-trivial to get largescale datasets of real DAGs to enable analysis and optimization. Taking workload execution as an example, DAGs capturing execution behavior can help engineers gain valuable insights into performance metrics such as latency and resource consumption exhibited by candidate systems (Luo et al., 2021; Cortez et al., 2017) to optimize the software and hardware accordingly. However, collecting a workload execution DAG (Sridharan et al., 2023) for even a single Large Language Model training job, potentially involving trillions of operations, is extremely prohibitive given the size of modern AI platforms. For example, Meta’s Llama3 was trained on a 24K GPU platform (Dubey et al., 2024). Moreover, very few companies in the world even have access to platforms of this scale. Even if such a large workload execution DAG could be collected, practical constraints such as the storage requirements of the DAGs, and potential to leak private information about the AI model architecture or the training platform configuration further limit DAG sharing. Furthermore, a representative and smaller graph is more effective for most performance optimization tasks.
This work focuses on developing generative models for DAGs, which allows various potential applications. First, being able to be data-driven and generate representative small set of DAGs can be much more compute and memory efficient than a real DAG given repetitive patterns. Second, this could enable data sharing without compromising intellectual property (Gao et al., 2024; Lin et al., 2020), thereby promoting software and hardware co-design among different companies (Sridharan et al., 2023). Third, a conditional DAG generative model can efficiently search the space of valid DAGs for optimization purposes in scenarios such as circuit design (Takagi, 1999; Dong et al., 2023), compiler optimization (Aho et al., 2006; Phothilimthana et al., 2023b), and neural architecture search (NAS) (Zhang et al., 2019; Li et al., 2023a; An et al., 2023).
Serving as an abstraction for flows and node dependencies, DAGs pose significant challenges in developing powerful and efficient generative models due to their intrinsic strong directional and logical dependencies, such as control flows, logic gates, and dimension requirements of matrix operations (as illustrated in Fig. 1 (a)). These complexities are further magnified in large-scale DAGs, presenting a unique combination of challenges regarding both scale and logical rules. In contrast, although social networks may be large in scale, they do not exhibit strong dependencies between nodes and edges. At the other end of the spectrum, molecular graphs, whose generation also has received much interest recently, must adhere to chemical rules but are generally smaller in size.
@@
Our model advances existing DAG generative models in multiple aspects (Zhang et al., 2019; Li et al., 2023a; An et al., 2023). Methodologically, although autoregressive models have been adopted by D-VAE (Zhang et al., 2019) and GraphPNAS (Li et al., 2023a) for DAG generation, they treat either a single node or a node set of constant size as a token. This tokenization method imposes an order between nodes that should be incomparable in the partial order, violating the inductive bias inherent in the DAG structure. We argue that the violation may hurt the generalization capability of generative models. Moreover, D-VAE suffers from a potentially slow generation process by generating one node per iteration. Although GraphPNAS generates multiple nodes per iteration, it uses a mixture of Bernoulli distributions to model intra-set edge dependencies, which is less expressive than diffusion models (Wang et al., 2023; Pearce et al., 2023). DiffusionNAG (An et al., 2023) employs diffusion models to generate only node attributes given a DAG structure. Other diffusion models have been developed for undirected graphs (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023a), but they ignore the directional information in DAGs, while our work demonstrates the necessity of the autoregressive component in modeling directional dependencies in DAGs.
@@ 2 Preliminaries and background
DAG generation A DAG is a directed graph without cycles $G=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\{1,\cdots,N\}$ is the node set, =$\mathcal{E}=\{\cdots,(u,v),\cdots\}$ is the set of directed edges, $\mathbf{X}$ is the node attribute matrix. The edge list $\mathcal{E}$ can also be equivalently represented as the adjacency matrix $\mathbf{A}$. From a collection of DAGs $\{G_{i}\}_{i}\sim\mathbb{P}_{\mathcal{G}}$, we aim to learn $\mathbb{P}_{\mathcal{G}}$ with a generative model $\mathbb{P}_{\mathcal{G}}^{\theta}$, where $\theta$ denotes model parameters. Additionally, many real-world applications, especially computing system benchmarking, involve labeled DAGs $\{(G_{i},y_{i})\}_{i}$ where the labels indicate the performance of the corresponding computing systems. DAG generation conditioning on these labels is also of interest.
Discrete denoising diffusion While our framework is compatible with various diffusion models, we adopt Discrete Denoising Diffusion Probabilistic Model (D3PM) (Austin et al., 2021) as a proof of concept. We briefly review it here. D3PM (Austin et al., 2021) generalizes Multinomial Diffusion (Hoogeboom et al., 2021b) in extending DDPM (Ho et al., 2020) to discrete data. D3PM has two phases, a forward diffusion process and a reverse process. Let each entry of $\mathbf{Z}^{(0)}\in\{0,1\}^{M\times C}$ be an one-hot encoding of a categorical attribute with $C$ possible values. The forward process uses $T$ consecutive steps to progressively corrupt $\mathbf{Z}^{(0)}\rightarrow\mathbf{Z}^{(1)}\rightarrow\cdots$ into purely random variables $\mathbf{Z}^{(T)}$ where $\mathbf{Z}^{(T)}$ can be drawn from a prior categorical distribution. To corrupt $\mathbf{Z}^{(t)}$ into $\mathbf{Z}^{(t+1)}$, it computes and samples from a conditional distribution $q(\mathbf{Z}^{(t+1)}|\mathbf{Z}^{(t)},t)=\mathbf{Z}^{(t)}\mathbf{Q}^{(t+1)}$, where $\mathbf{Q}^{(t+1)}\in\mathbb{R}^{C\times C}$ is a pre-determined transition matrix. A denoising network $\phi_{\theta}$ is trained to predict the uncorrupted data $\mathbf{Z}^{(0)}$ from $(\mathbf{Z}^{(t)},t)$. During the reverse generation process, the trained denoising network is used to convert $\mathbf{Z}^{(T)}$ drawn from the prior distribution into realistic data. DiGress (Vignac et al., 2023a) extends D3PM for undirected molecular graph generation by treating the presence or absence of an edge between a node pair as a binary attribute. The denoising of node attributes and edges takes the form of node classification and link prediction, respectively.
@@ 3.1 A unique sequential view of DAGs
The structure of a DAG essentially establishes a partial order among its nodes based on reachability through directed paths. A partially ordered set can be transformed into an ordered sequence while preserving compatibility with the partial order, known as a linear extension or a topological ordering in the context of DAGs (Skiena, 2008). Previous competitive autoregressive models of DAGs utilize this sequential nature to generate nodes one at a time following a topological ordering (Zhang et al., 2019). However, topological orderings are not unique. This may compromise the model’s effectiveness and efficiency. We leave a dedicated discussion about the issue to Section 3.3. On the other hand, non-autoregressive models like diffusion models may be sub-optimal for not explicitly modeling the sequential nature of DAGs.
@@ 3.2 LayerDAG
This factorization naturally leads to a layerwise autoregressive generation framework. Iteratively, to generate the $(l+1)^{\text{th}}$ layer, it first predicts the number of new nodes with $p_{\theta}\left(|\mathcal{V}^{(l+1)}|\mid G^{(\leq l)}\right)$. Then it generates the node attributes with $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$. Finally, it generates the edges with $p_{\theta}\left(\mathbf{A}^{(l+1)}\mid G^{(\leq l)},\mathbf{X}^{(l+1)}\right)$. The generation process terminates when $|\mathcal{V}^{(l+1)}|$ is predicted to be 0. With layerwise autoregressive generation, this framework potentially allows better generalization to unseen values of $L$ during generation (layer generalization).
@@
Layerwise diffusion-based generation To preserve the uniqueness of the layerwise tokenization, we need to generate the set of node attributes and edges in the bipartite graph $G^{(l+1)}$ as a whole by modeling $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$ and $p_{\theta}\left(\mathbf{A}^{(l+1)}\mid G^{(\leq l)},\mathbf{X}^{(l+1)}\right)$ as set-level generations. This requires capturing complex dependencies between the nodes attributes and edges in the sets, which is crucial for the applications like system and circuit design, where strict logical rules are prevalent, and rule violations can accumulate and propagate over layers of generation. To tackle this issue, we adopt diffusion models for multiple rounds of set refinement in the generation of $G^{(l+1)}$. We employ two separate diffusion processes for node attribute $\mathbf{X}^{(l+1)}$ generation and structure $\mathbf{A}^{(l+1)}$ generation, as suggested by Li et al. (2023b). For simplicity, we focus on categorical node attributes in this work, which find abundant real-world applications like operator types in computational graphs, and thus adopt D3PM (Austin et al., 2021). Our approach can be extended to handle real-valued attributes by combining discrete diffusion with continuous diffusion, as demonstrated in previous works (Vignac et al., 2023b; Hua et al., 2024). For more details on layerwise diffusion, see Appendix A.
@@
Implementation LayerDAG consists of three separate modules $p_{\theta}\left(|\mathcal{V}^{(l+1)}|\mid G^{(\leq l)}\right)$, $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$, and $p_{\theta}\left(\mathbf{A}^{(l+1)}\mid G^{(\leq l)},\mathbf{X}^{(l+1)}\right)$, whose parameters are shared across layers. All three modules involve a DAG encoder. We use an off-the-shelf bidirectional message passing neural network (BiMPNN) (Wen et al., 2020), which has been proven effective for encoding directed graphs in hardware systems (Wang et al., 2024). A single BiMPNN layer updates node representations with synchronous message passing over both the directed edges and their reversed counterparts: $\sigma\left(\mathbf{A}\mathbf{H}\mathbf{W}_{\text{forward}}+\mathbf{A}^{\top}\mathbf{H}\mathbf{W}_{\text{reverse}}+\mathbf{H}\mathbf{W}_{\text{self}}\right)$, where $\sigma$ is a non-linear layer, $\mathbf{H}$ is the node representation matrix, and $\mathbf{W}$’s are learnable weight matrices. Both layer size generation $p_{\theta}\left(|\mathcal{V}^{(l+1)}|\mid G^{(\leq l)}\right)$ and node attribute generation $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$ involve computing graph representations with a set pooling operator over the updated node representations. Let $\mathbf{X}^{(l+1,t)}$ be the node attributes sampled for the $(l+1)^{\text{th}}$ layer at the step $t$ in the reverse node attribute diffusion process. After encoding $(G^{(\leq l)},t)$ into a context vector $\mathbf{h}_{t}^{(\leq l)}$, the denoising network $\phi_{\theta_{X}}$ augments the representations of $\mathbf{X}^{(l+1,t)}$ by $\mathbf{h}_{t}^{(\leq l)}$ and then applies a transformer without positional encodings (Vaswani et al., 2017) over them for predicting the set of $\mathbf{X}^{(l+1)}$. Similarly, the denoising network $\phi_{\theta_{A}}$ for edge diffusion augments $(G^{(\leq l)},t)$ by $(\mathbf{X}^{(l+1)},\mathbf{A}^{(l+1,t)})$ for computing the node representations of $\mathcal{V}^{(\leq l+1)}$. To predict the probability of edge $(u,v)$ for $u\in\mathcal{V}^{(\leq l)}$ and $v\in\mathcal{V}^{(l+1)}$, it concatenates and transforms the representations of node $u$, node $v$, and $t$ with an MLP.
@@
Layerwise positional encoding An important out-of-distribution generalization issue for sequence generation is length generalization, which challenges language models trained on shorter sequences to extrapolate to longer ones during generation (Anil et al., 2022; Zhou et al., 2024). A substantial body of research approaches this issue by developing more generalizable positional encodings (PEs) for tokens (Shaw et al., 2018; Su et al., 2024; Ruoss et al., 2023).
As DAGs generalize chains/sequences, we consider the use of PEs based on the layer index $l$ to enhance model expressive power. In our evaluation, we notice that PEs based on one hot encoding of the layer index $l$ may hurt the performance. This observation is different from D-VAE (Zhang et al., 2019) that uses one-hot encodings of node indices given by a topological order. Relative PEs based on index difference between the latest layer and existing layers, are translation invariant and compatible with layer generalization. However, they are not expressive enough to capture layer-specific patterns. Therefore, we propose to use the sinusoidal encoding of $l$ given by $(\cdots,\cos(w_{i}l),\sin(w_{i}l),\cdots)$ as PEs for nodes in layer $l$, which improves the expressive power while keeping a certain level of generalization. This choice of PE improves the final quality of the generated DAGs in many cases.

### v2

@@ 1 Introduction
A Directed Acyclic Graph (DAG) is a data structure that represents the order of elements (Bang-Jensen & Gutin, 2008). Unlike linear sequences, which follow a single, straight path from the first to the last element, DAGs incorporate branching and merging, allowing for the modeling of complex dependencies and hierarchies. This flexibility makes DAGs ideal for representing diverse problem domains such as workload behavior during system execution Sridharan et al. (2023), operator dependence for program analysis (Phothilimthana et al., 2023b; Luo et al., 2021; Cortez et al., 2017), dataflows in circuits (Dong et al., 2023), task dependencies in project management (Skiena, 2008; Borenstein, 2000), and cause-effect relationships (Pearl, 1995; Tennant et al., 2020). Additionally, DAGs have recently been used to create challenging benchmarks for evaluating the reasoning capabilities of large language models (Zhu et al., 2024).
@@
Despite the benefits of DAGs mentioned above, it is non-trivial to get largescale datasets of real DAGs to enable analysis and optimization. Taking workload execution as an example, DAGs capturing execution behavior can help engineers gain valuable insights into performance metrics such as latency and resource consumption exhibited by candidate systems (Luo et al., 2021; Cortez et al., 2017) to optimize the software and hardware accordingly. However, collecting a workload execution DAG (Sridharan et al., 2023) for even a single Large Language Model training job, potentially involving trillions of operations, is extremely prohibitive given the size of modern AI platforms. For example, Meta’s Llama3 was trained on a 24K GPU platform (Dubey et al., 2024) while xAI is building a 100K GPU platform. Moreover, few companies in the world even have access to platforms of this scale. Even if such a large workload execution DAG could be collected, practical constraints such as the storage requirements of the DAGs, and potential to leak private information about the AI model architecture or the training platform configuration further limits DAG sharing. Furthermore, a representative and smaller graph is more effective for most performance optimization tasks.
This work focuses on developing generative models for DAGs. We believe this can solve several challenges. First, being able to be data-driven and generate representative small set of DAGs can be much more compute and memory efficient than a real DAG given repetitive patterns. Second, this could enable data sharing without compromising intellectual property (Gao et al., 2024; Lin et al., 2020), thereby promoting software and hardware co-design among different companies (Sridharan et al., 2023). Third, a conditional DAG generative model can efficiently search the space of valid DAGs for optimization purposes in scenarios such as circuit design (Takagi, 1999; Dong et al., 2023), compiler optimization (Aho et al., 2006; Phothilimthana et al., 2023b), and neural architecture search (NAS) (Zhang et al., 2019; Li et al., 2023a; An et al., 2023).
Serving as an abstraction for flows and node dependencies, DAGs pose significant challenges in developing powerful and efficient generative models due to their intrinsic strong directional and logical dependencies, such as control flows, logic gates, and dimension requirements of matrix operations (as illustrated in Fig. 1 (a)). These complexities are further magnified in large-scale DAGs, presenting a unique combination of challenges regarding both scale and logical rules. In contrast, although social networks may be large in scale, they do not exhibit strong dependencies between nodes and edges. At the other end of the spectrum, molecular graphs, whose generation has also received much interest recently, must adhere to chemical rules but are generally smaller in size.
@@
Our model advances existing DAG generative models in multiple aspects (Zhang et al., 2019; Li et al., 2023a; An et al., 2023). Methodologically, although autoregressive models have been adopted by D-VAE (Zhang et al., 2019) and GraphPNAS (Li et al., 2023a) for DAG generation, they treat either a single node or a node set of constant size as a token. This tokenization method imposes an order between nodes that should be incomparable in the partial order, violating the inductive bias inherent in the DAG structure. We argue that the violation may hurt the generalization capability of generative models. Moreover, D-VAE suffers from a potentially slow generation process by generating one node per iteration. Although GraphPNAS generates multiple nodes per iteration, it uses a mixture of Bernoulli distributions to model intra-set edge dependencies, which is less expressive than diffusion models (Wang et al., 2023; Pearce et al., 2023). DiffusionNAG (An et al., 2023) employs diffusion models to generate only node attributes given a DAG structure. Diffusion models have been used to generate undirected graphs (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023a), but they ignore the directional information in DAGs, while our work demonstrates the necessity of the autoregressive component in modeling directional dependencies in DAGs.
@@ 2 Preliminaries and background
DAG generation A DAG is a directed graph without cycles $G=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\{1,\cdots,N\}$ is the node set, =$\mathcal{E}=\{\cdots,(u,v),\cdots\}$ is the set of directed edges, $\mathbf{X}$ is the node attribute matrix. The edge list $\mathcal{E}$ can also be equivalently represented as the adjacency matrix $\mathbf{A}$. From a collection of DAGs $\{G_{i}\}_{i}\sim\mathbb{P}_{\mathcal{G}}$, we aim to learn $\mathbb{P}_{\mathcal{G}}$ with a generative model $\mathbb{P}_{\mathcal{G}}^{\theta}$, where $\theta$ denotes model parameters. Additionally, many real-world applications, especially computing system benchmarking, involve labeled DAGs $\{(G_{i},y_{i})\}_{i}$ where the labels indicate the performance of the corresponding computing systems. DAG generation conditioning on these labels is also of interest, i.e., learning the distribution $\mathbb{P}_{\mathcal{G}|\mathcal{Y}}$.
Discrete denoising diffusion While our framework is compatible with various diffusion models, we adopt Discrete Denoising Diffusion Probabilistic Model (D3PM) (Austin et al., 2021) as a proof of concept. We briefly review it here. D3PM (Austin et al., 2021) generalizes Multinomial Diffusion (Hoogeboom et al., 2021b) in extending DDPM (Ho et al., 2020) to discrete data. D3PM has two phases, a forward diffusion process and a reverse process. Let each entry of $\mathbf{Z}^{(0)}\in\{0,1\}^{M\times C}$ be a one-hot encoding of a categorical attribute with $C$ possible values.
The forward process uses $T$ consecutive steps to progressively corrupt $\mathbf{Z}^{(0)}\rightarrow\mathbf{Z}^{(1)}\rightarrow\cdots$ into purely random variables $\mathbf{Z}^{(T)}$ where $\mathbf{Z}^{(T)}$ can be drawn from a prior categorical distribution. To corrupt $\mathbf{Z}^{(t)}$ into $\mathbf{Z}^{(t+1)}$, it computes and samples from a conditional distribution $q(\mathbf{Z}^{(t+1)}|\mathbf{Z}^{(t)},t)=\mathbf{Z}^{(t)}\mathbf{Q}^{(t+1)}$, where $\mathbf{Q}^{(t+1)}\in\mathbb{R}^{C\times C}$ is a pre-determined transition matrix. A denoising network $\phi_{\theta}$ is trained to predict the uncorrupted data $\mathbf{Z}^{(0)}$ from $(\mathbf{Z}^{(t)},t)$. During the reverse generation process, the trained denoising network is used to convert $\mathbf{Z}^{(T)}$ drawn from the prior distribution into realistic data. DiGress (Vignac et al., 2023a) extends D3PM for molecular graph (undirected) generation by treating the presence or absence of an edge between a node pair as a binary attribute. The denoising of node attributes and edges takes the form of node classification and link prediction, respectively.
@@ 3.1 A unique sequential view of DAGs
The structure of a DAG essentially establishes a partial order among its nodes based on reachability through directed paths. A partially ordered set can be transformed into an ordered sequence while preserving compatibility with the partial order, known as a linear extension or a topological ordering in the context of DAGs (Kahn, 1962; Tarjan, 1972). Previous competitive autoregressive models of DAGs utilize this sequential nature to generate nodes one at a time following a topological ordering (Zhang et al., 2019). However, topological orderings are not unique. This may compromise the model’s effectiveness and efficiency. We leave a dedicated discussion about the issue to Section 3.3. On the other hand, non-autoregressive models like diffusion models may be sub-optimal for not explicitly modeling the sequential nature of DAGs.
@@ 3.2 LayerDAG
This factorization naturally leads to a layerwise autoregressive generation framework. Iteratively, to generate the $(l+1)^{\text{th}}$ layer, it first predicts the number of new nodes with $p_{\theta}\left(|\mathcal{V}^{(l+1)}|\mid G^{(\leq l)}\right)$. Then it generates the node attributes with $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$. Finally, it generates the edges with $p_{\theta}\left(\mathbf{A}^{(l+1)}\mid G^{(\leq l)},\mathbf{X}^{(l+1)}\right)$. The generation process terminates when $|\mathcal{V}^{(l+1)}|$ is predicted to be 0. All the above generation processes will be conditioned on the previously generated DAGs $G^{(\leq l)}$. With layerwise autoregressive generation, this framework potentially allows better generalization to unseen values of $L$ during generation (layer generalization).
@@
Layerwise diffusion-based generation To preserve the uniqueness of the layerwise tokenization, we need to generate the set of node attributes and edges in the bipartite graph $G^{(l+1)}$ as a whole by modeling $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$ and $p_{\theta}\left(\mathbf{A}^{(l+1)}\mid G^{(\leq l)},\mathbf{X}^{(l+1)}\right)$ as set-level generations. This requires capturing complex dependencies between the node attributes and edges in the sets, which is crucial for applications like system and circuit design, where strict logical rules are prevalent, and rule violations can accumulate and propagate over layers of generation. To tackle this issue, we adopt diffusion models for multiple rounds of set refinement in the generation of $G^{(l+1)}$. We employ two separate diffusion processes for node attribute $\mathbf{X}^{(l+1)}$ generation and structure $\mathbf{A}^{(l+1)}$ generation, as suggested by Li et al. (Li et al., 2023b). For simplicity, we focus on categorical node attributes in this work, which find abundant real-world applications like operator types in computational graphs, and thus adopt D3PM (Austin et al., 2021). Our approach can be extended to handle real-valued attributes by combining discrete diffusion with continuous diffusion, as demonstrated in previous works (Vignac et al., 2023b; Hua et al., 2024). For more details on layerwise diffusion, see Appendix A.
@@
Implementation For different $l$’s, all modules $p_{\theta}\left(|\mathcal{V}^{(l+1)}|\mid G^{(\leq l)}\right)$, $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$, and $p_{\theta}\left(\mathbf{A}^{(l+1)}\mid G^{(\leq l)},\mathbf{X}^{(l+1)}\right)$ share the same parameter $\theta$, which involves a DAG encoder. We use an off-the-shelf bidirectional message passing neural network (BiMPNN) (Wen et al., 2020). A single BiMPNN layer updates node representations with synchronous message passing over both the directed edges and their reversed counterparts: $\sigma\left(\mathbf{A}\mathbf{H}\mathbf{W}_{\text{forward}}+\mathbf{A}^{\top}\mathbf{H}\mathbf{W}_{\text{reverse}}+\mathbf{H}\mathbf{W}_{\text{self}}\right)$, where $\sigma$ is a non-linear layer, $\mathbf{H}$ is the node representation matrix, and $\mathbf{W}$’s are learnable weight matrices. Both layer size generation $p_{\theta}\left(|\mathcal{V}^{(l+1)}|\mid G^{(\leq l)}\right)$ and node attribute generation $p_{\theta}\left(\mathbf{X}^{(l+1)}\mid G^{(\leq l)},|\mathcal{V}^{(l+1)}|\right)$ involve computing graph representations with a set pooling operator over the updated node representations. Let $\mathbf{X}^{(l+1,t)}$ be the node attributes sampled for the $(l+1)^{\text{th}}$ layer at the step $t$ in the reverse node attribute diffusion process. After encoding $(G^{(\leq l)},t)$ into a context vector $\mathbf{h}_{t}^{(\leq l)}$, the denoising network $\phi_{\theta_{X}}$ augments the representations of $\mathbf{X}^{(l+1,t)}$ by $\mathbf{h}_{t}^{(\leq l)}$ and then applies a transformer without positional encodings (Vaswani et al., 2017) over them for predicting the set of $\mathbf{X}^{(l+1)}$. Similarly, the denoising network $\phi_{\theta_{A}}$ for edge diffusion augments $(G^{(\leq l)},t)$ by $(\mathbf{X}^{(l+1)},\mathbf{A}^{(l+1,t)})$ for computing the node representations of $\mathcal{V}^{(\leq l+1)}$. To predict the probability of edge $(u,v)$ for $u\in\mathcal{V}^{(\leq l)}$ and $v\in\mathcal{V}^{(l+1)}$, it concatenates and transforms the representations of node $u$, node $v$, and $t$ with an MLP.